OpenStudy (anonymous):

Find the equation for the cubic function, f(x), with roots at -5, 5 and -3, and has a y-intercept at (0, 2).

OpenStudy (anonymous):

x^3+3x^2-25x+2 The concept here is that knowing the roots allow you to formulate the terms y=(x+5)(x-5)(x+3) however, you still need a constant term C in this equation y=(x+5)(x-5)(x+3)+C and you can find C by saying (x,y) =(0,2) the constant C is not obvious, but you should know that it exists because when you plug in (0,2) into the y function without C, your function is invalid thus, after you find C=77 expand the x and you will get y=(x^2-25)(x+3)+77 y=x^3+3x^2-25x-75+77 y=x^3+3x^2-25x+2, which is your answer! GOOD JOB!

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

:D of course